OpenStudy (i_ask_alot):

I need help with 5 questions. 1.A jar contains 3 oz of seed and is 1/6 full. How much seed is in a full jar? 2.Alan can chop 3/4 of a carrot in 1/8 min. How many carrots can he chop in 1 min? 3.A clerk takes 3/4 h to process 9 applications. How long does it take the clerk to process 1 application? 4.A bulldozer moves 640 lb of dirt in 4/5 h. Complete the unit rate. 5.A contractor takes 4 1/2 days to build 3/4 of a deck. Complete the unit rate.

OpenStudy (anonymous):

1) It contains 3 oz, and is 1/6 full. So when it is full it contains \[3*6=?\]

OpenStudy (i_ask_alot):

@Tommynaiter is 18oz the answer to #1

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (i_ask_alot):

@ShiBuringa @Lexi_Loves Help???

OpenStudy (anonymous):

2) She can chop 3/4 in 1/8 min. So we need to find how long it takes to chop a full carrot. So you need to multiply the time with 4/3 because: \[\frac{3}{4}*\frac{4}{3}=1\] So you get: \[\frac{1}{8}*\frac{4}{3}=?\]

OpenStudy (anonymous):

3) If it takes 3/4 h, to do 9 applications, then you need to divide the time with 9 to get the time it takes to do 1 application. \[\frac{3/4}{9}=?\]

OpenStudy (i_ask_alot):

So is #2 1/6

OpenStudy (i_ask_alot):

And #3 1/12

OpenStudy (i_ask_alot):

@Tommynaiter

OpenStudy (i_ask_alot):

@iGreen

OpenStudy (anonymous):

2) Is correct, 1/6. 3) is correct, 1/12

OpenStudy (i_ask_alot):

Thanks

OpenStudy (i_ask_alot):

What about 4 and 5 do you mind helping with those

OpenStudy (anonymous):

4) It moves 640 lb in 4/5 h. How much does it move in 1 hour? You need to find how much you have to multiply 4/5 to get 1. This is just like 2).

OpenStudy (anonymous):

\[\frac{4}{5}*\frac{5}{4}=1\] Now multiply 640 with 5/4

OpenStudy (anonymous):

It is the same for 5), figure out how much you need to multiply to get 1.